Two transverse sinusoidal waves combining in a string are described by the wave functions y1 = 0.02 sin⁡(2πx+πt) and y2 = 0.02 sin(2πx-πt), where x and y are in meters, and t is in seconds. If after superposition, a standing wave of 3 loops is formed, then the length of the string is:


2 m


1.5 m


0.75 m


0.5 m


1 m


1
Expert's answer
2021-05-30T13:26:16-0400

"y_1=0.02\sin(2\pi x+\pi t)"

"y_2=0.02\sin(2\pi x-\pi t)"


Resultant wave, "y=y_1+y_2"

"y=0.04\sin(2\pi x)\cos(\pi t)"

Comparing this with standard equation,

"y=A\sin(kx)\cos(\omega t)"

"k=2\pi"

Also,

"k=2\pi\lambda"

"\lambda=1"


Number of loops, "n=3"

Length of string = "L"

Number of loops, "n=2L\lambda"

"L=\dfrac{n}{2\lambda}"

"L=\dfrac{3}{2}=1.5\space m"


Length of string = 1.5 m